# Extraction of Singular Patterns from a Vector Field via Persistent Path Homology

### v2

@@ VI-A Experiment Results of Extracting Singular Patterns
We also conducted another experiment in which we used projected geomagnetic field data to determine the location of the magnetic dip poles. The magnetic dip poles are defined as the locations where the main magnetic field as a whole is normal to Earth’s surface , which provide a crucial reference for local orientation when navigating on or close to Earth’s surface at high-latitudes [38]. However, observations have shown that the positions of the two magnetic dip poles change from year to year [38], so it is necessary to calculate and update their positions in time. Our approach is well-suited to address this issue. When we project geomagnetic field data onto a map, resulting in a planar vector field, the two dip poles (north dip pole and south dip pole) become singularities of this planar vector field, since the projected vectors are zero at these points. In this experiment, we select two domains around the two dip poles respectively and choose geomagnetic field data at zero altitude. After projecting onto the map, discrete vector fields sampled around the two dip poles are obtained. Using our method, we determine the locations of the two dip poles for the years 2000, 2005, 2010, 2015, and 2020, and compare our results with data of actual location provided by IGRF [38], as shown in Table II. It is evident that our method incurs only minor positioning errors, thus reaffirming its effectiveness. The positioning results are also plotted on maps (Fig. 11).
Figure 11: The positioning results of north dip pole (left) and south dip pole (right) for the years 2000, 2005, 2010, 2015, and 2020.
VI-B Vector Field Difference Measurement
In the following, we will use the wind field datasets as an example to demonstrate how our method can measure the variation of time-varying vector fields. Similar as the work [50], we will use the resulting persistence diagrams obtained by wind fields to measure the variations of singular patterns of the wind fields of the two tropical cyclones respectively.
From blue line in Fig. 12, it is evident that the bottleneck distance reached a local minimum at the moment 2023-08-03-18:00 UTC, coinciding with the time when Khanun became stationary near a position since a notable change in trajectory was taking place. This period exhibited reduced variability in the wind field, resulting in a smaller difference between the persistence diagrams. Similarly, in orange line in Fig. 12, at the moment 2023-08-27-18:00 UTC, Saola became stationary near a position, where its trajectory was changing direction, indicating small variability between the persistence diagrams of the corresponding wind fields. Similar features of bottleneck distance can be observed at other moments, such as 2023-08-31-06:00 UTC, during which period Saola moves slowly within the vicinity of a location. Conversely, if a tropical cyclone undergoes significant movement or multiple turns over a period of time, the corresponding bottleneck distance will be larger. Moreover, a significant fluctuation of the bottleneck distance may indicate that the tropical cyclone is undergoing unstable movement and change.
VI-C Comparisons with State-of-the-art Methods
In this section, we will compare the mean positioning errors and running time between our method and several state-of-the-art methods. To calculate the mean errors, suppose that we have data for $n$ moments, and we denote the $i$-th actual longitude as $al_{i}$ and the $i$-th positioning longitude as $pl_{i}$. The mean longitude error can be computed using the formula $\frac{1}{n}\sum_{i=1}^{n}\left|al_{i}-pl_{i}\right|$, and the mean latitude error can be computed similarly. Table III provide a comparison of the mean errors obtained using our method with some state-of-the-art singularity extraction methods, they are methods based on Hodge decomposition [24], triangular linear interpolation of vectors [25], and computation of the Jacobian matrix [26]. The average running times are also shown in Table III. The Khanun datasets consists of 1600 sampling vectors (resulting in a digraph of 1600 vertices and 3120 edges), the Saola datasets comprises 1924 sampling vectors (resulting in a digraph of 1924 vertices and 3759 edges), and the geomagnetic field datasets consists of 1600 sampling vectors (resulting in a digraph of 1600 vertices and 3120 edges). Based on the results, our method achieves smaller errors than the method proposed by [24], and its time cost is much shorter than [24]. Furthermore, our approach can achieve errors nearly as small as the numerical methods of [25] and [26], although it has a slightly longer time cost than those numerical methods. However, these numerical methods may fail to correctly identify singularities or encounter spurious singularities due to precision issues. For instance, in the experiments, both the triangular linear interpolation method and the Jacobian matrix method yield spurious singularities. The number of spurious singularities obtained from all vector fields in the corresponding datasets is shown in Table III. For example, in the geomagnetic field datasets, there are 10 vector fields and 10 corresponding singularities in total, but the triangular interpolation method finds 15 singularities, including 5 spurious singularities, and the Jacobian method finds 12 singularities, including 2 spurious singularities. In contrast, our method, which relies on the singularity index, finds all of the singularities without any spurious one. Additionally, our method has the capability to measure the variations between vector fields, while the other methods are unable to accomplish this.
Figure 12: Bottleneck distances computed by one-dimensional persistence diagrams derived from Khanun and Saola datasets respectively.
TABLE III: Comparative Results
Comparative Results for Khanun Datasets (36 vector fields)
Triangular interpolation [25] 0.134 0.107 27.21 15
Jacobian method [26] 0.126 0.109 45.25 9
Hodge decomposition [24] 0.205 0.147 566.64 0
our method 0.131 0.118 70.19 0
Comparative Results for Saola Datasets (32 vector fields)
Triangular interpolation [25] 0.108 0.093 34.76 15
Jacobian method [26] 0.111 0.093 56.00 2
Hodge decomposition [24] 0.129 0.133 689.39 0
our method 0.110 0.102 103.08 0
Triangular interpolation [25] 0.028 0.056 27.90 5
Jacobian method [26] 0.030 0.000 47.31 2
Hodge decomposition [24] 0.089 0.182 572.02 0
our method 0.037 0.059 90.95 0
VI-D Discussion
Our method exhibits several limitations. If the sampling grid is not sufficiently dense, some singularities may not be extracted because multiple singularities may exist within the same square, and their indices can cancel each other out. Furthermore, singularities located on the edges of the grid digraph may be overlooked. In general, these situations are unlikely to occur due to practical considerations and limitations in numerical accuracy. However, new techniques should be developed to address these theoretical issues. Furthermore, the singular polygon may offer a coarse approximation of the singular pattern, given that the minimal cells in grid digraphs are square, implying that the boundary may not be smooth. Additionally, the existence of noise may cause a few edges to exist within the polygon, contradicting Theorem 5. For instance, in Fig. 10, even edges on the same line on the singular polygon may have different directions, although the proportion of such edges is minimal. Moreover, if the singularity and its corresponding singular pattern are located too close to the boundary of the considered region, the extraction of the singular polygon may fail. Therefore, further refinement is required to obtain a more accurate representation of the region exhibiting the singular pattern.
VII Conclusions
In this study, we propose a novel method for determining the position of a singularity and identifying the corresponding region of the singular pattern (referred to as the singular polygon) within a planar vector field exhibiting a singular pattern. Moreover, we compare the changes of singular patterns of time-varying vector fields using persistence diagrams. Our approach involves the conversion of discrete vector fields into angle-based grid digraphs, which are then analyzed using persistent path homology. Experimental results demonstrate that our proposed method achieves low mean errors in locating the centers of tropical cyclones and locations of dip poles, and effectively measures the changes in tropical cyclone features over time. Additionally, we can obtain the regions where singular patterns are present.
In future work, we aim to explore alternative techniques for converting vector fields into digraphs to accommodate diverse applications, as well as more precise methods for extracting singular patterns. Furthermore, we will address challenges related to vector fields in higher dimensional Euclidean spaces or surfaces and investigate the potential utilization of higher dimensional (persistent) path homology as a new analytical tool.
Acknowledgments

### v3

@@ Appendix B Some tables of supplementary data
Table 3: Positioning Results and Errors of the Center Tracking of Typhoon Saola
[…]
2023/08/25/00:00 124 19.7 124.26 19.99 0.26 0.29
2023/08/25/06:00 123.9 19.7 124.25 19.52 0.35 0.18
2023/08/25/12:00 123.6 19.5 123.96 19.26 0.36 0.24
2023/08/25/18:00 123.4 19.3 123.53 19.22 0.13 0.08
2023/08/26/00:00 123.2 18.5 123.26 18.71 0.06 0.21
2023/08/26/06:00 123.2 18 123.28 18 0.08 0
2023/08/26/12:00 123.1 17.6 123.02 17.52 0.08 0.08
2023/08/26/18:00 122.9 17.2 123.04 17.24 0.14 0.04
2023/08/27/00:00 122.9 16.8 123 16.77 0.1 0.03
2023/08/27/06:00 123 16.5 122.97 16.49 0.03 0.01
2023/08/27/12:00 123.2 16.3 123.28 16.23 0.08 0.07
2023/08/27/18:00 123.8 16 123.74 16.21 0.06 0.21
2023/08/28/00:00 124.3 16.8 124.27 16.51 0.03 0.29
2023/08/28/06:00 124.3 17.5 124.26 17.48 0.04 0.02
2023/08/28/12:00 124.1 17.8 124.25 17.75 0.15 0.05
2023/08/28/18:00 123.9 18.2 124 18.22 0.1 0.02
2023/08/29/00:00 123.5 18.6 123.51 18.51 0.01 0.09
2023/08/29/06:00 123.1 18.9 123 18.76 0.1 0.14
2023/08/29/12:00 122.7 19.3 122.51 19.29 0.19 0.01
2023/08/29/18:00 121.9 19.9 121.96 19.73 0.06 0.17
2023/08/30/00:00 121.1 20.1 121.23 20 0.13 0.1
2023/08/30/06:00 120.4 20.4 120.46 20.49 0.06 0.09
2023/08/30/12:00 119.7 20.7 119.51 20.71 0.19 0.01
2023/08/30/18:00 118.7 20.9 118.99 20.96 0.29 0.06
2023/08/31/00:00 118.2 21 118.22 21 0.02 0
2023/08/31/06:00 117.8 21.2 117.74 21.26 0.06 0.06
2023/08/31/12:00 117.3 21.5 117.25 21.47 0.05 0.03
2023/08/31/18:00 116.8 21.7 116.99 21.54 0.19 0.16
2023/09/01/00:00 116.3 21.9 116.22 21.97 0.08 0.07
2023/09/01/06:00 115.5 22 115.5 21.78 0 0.22
2023/09/01/12:00 114.5 22 114.52 22.01 0.02 0.01
2023/09/01/18:00 113.5 22 113.49 21.77 0.01 0.23
Table 4: Positioning Results and Errors of the Center Tracking of Typhoon Khanun
[…]
2023/07/31/00:00 132 22.1 131.99 22.22 0.01 0.12
2023/07/31/06:00 131.5 22.8 131.52 23 0.02 0.2
2023/07/31/12:00 131.1 23.4 130.99 23.29 0.11 0.11
2023/07/31/18:00 130.3 24 130.03 24.25 0.27 0.25
2023/08/01/00:00 129.4 24.6 129.27 24.51 0.13 0.09
2023/08/01/06:00 128.7 25 128.52 24.97 0.18 0.03
2023/08/01/12:00 128 25.3 127.77 25.22 0.23 0.08
2023/08/01/18:00 127.4 25.6 127.26 25.5 0.14 0.1
2023/08/02/00:00 126.8 25.7 126.78 25.75 0.02 0.05
2023/08/02/06:00 126.1 26 126 26.03 0.1 0.03
2023/08/02/12:00 125.6 26.2 125.52 26.02 0.08 0.18
2023/08/02/18:00 125.1 26.4 125.03 26.28 0.07 0.12
2023/08/03/00:00 124.8 26.7 124.77 26.53 0.03 0.17
2023/08/03/06:00 124.3 26.7 124.3 26.74 0 0.04
2023/08/03/12:00 124.1 26.9 124.26 26.72 0.16 0.18
2023/08/03/18:00 124.2 26.8 124.26 26.78 0.06 0.02
2023/08/04/00:00 124.5 27 124.72 27.02 0.22 0.02
2023/08/04/06:00 125.2 27.2 125.21 27.25 0.01 0.05
2023/08/04/12:00 125.7 27.5 125.73 27.49 0.03 0.01
2023/08/04/18:00 126.3 27.6 126.23 27.55 0.07 0.05
2023/08/05/00:00 126.9 27.7 126.75 27.77 0.15 0.07
2023/08/05/06:00 127.6 27.8 127.75 27.97 0.15 0.17
2023/08/05/12:00 128.4 27.9 128.47 27.78 0.07 0.12
2023/08/05/18:00 129 27.9 128.96 28.22 0.04 0.32
2023/08/06/00:00 129.4 27.8 129.5 28.2 0.1 0.4
2023/08/06/06:00 130 27.7 129.98 27.51 0.02 0.19
2023/08/06/12:00 130.5 27.5 130.3 27.46 0.2 0.04
2023/08/06/18:00 130.8 27.7 130.77 27.72 0.03 0.02
2023/08/07/00:00 131 27.8 131.25 27.75 0.25 0.05
2023/08/07/06:00 131.1 27.9 131.26 27.98 0.16 0.08
2023/08/07/12:00 131.3 28.2 131 28.24 0.3 0.04
2023/08/07/18:00 131.1 28.4 130.75 28.54 0.35 0.14
2023/08/08/00:00 130.9 29 131.04 29.24 0.14 0.24
2023/08/08/06:00 130.8 29.3 130.78 29.53 0.02 0.23
2023/08/08/12:00 130.4 29.9 130 29.78 0.4 0.12
2023/08/08/18:00 129.7 30.4 129.29 30.53 0.41 0.13